OpenStudy (anonymous):

Which of the following angles is in the solution set of csc^2θ - cscθ - 20 = 0 for 0° ≤ θ < 360° ans:14.5°,165.5°,168.5°,348.5°

OpenStudy (mertsj):

Factor out csc theta Can you do that?

OpenStudy (mertsj):

Oh> Just a minute. Your post isn't clear. Why don't you retype it and use x for the angle. Then it might be clearer.

OpenStudy (mertsj):

is it really -20?

OpenStudy (mertsj):

\[\csc ^2\theta-\csc \theta-20=0\]

OpenStudy (anonymous):

yes it is

OpenStudy (mertsj):

Is that it?

OpenStudy (mertsj):

\[x^2-x-20=0\]

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Which has these factors: \[(x-5)(x+4)=0\]

OpenStudy (mertsj):

So we have: \[(\csc \theta-5)(\csc \theta+4)=0\]

OpenStudy (mertsj):

\[\csc \theta=5 and \csc \theta=-4\]

OpenStudy (anonymous):

i put the answer options under the question.

OpenStudy (mertsj):

so \[\sin \theta=\frac{1}{5}\]

OpenStudy (mertsj):

and \[\sin \theta=-\frac{1}{4}\]

OpenStudy (mertsj):

So put those values into your calculator and hit the inverse sin button

OpenStudy (anonymous):

it doesnt come out to any of the answers

OpenStudy (mertsj):

What did you get when you put in -1/4?

OpenStudy (anonymous):

the sin of -1/4 is -0.004363309 and the inverse is -0.25

OpenStudy (mertsj):

Find \[\sin ^{-1}(.2)\]

OpenStudy (mertsj):

Then find \[\sin ^{-1}(-.25)\]

OpenStudy (anonymous):

i got 11.53695903 and then -14.47751219

OpenStudy (anonymous):

in radian mode i got 0.201357921 and -0.252680255 the other was in degree mode